An amount of 40% acid solution is to be mixed with enough 10% acid solution to make 25% acid solution. If there are to be 20 litres of the final mixture, how much of each solution should be mixed together?
Answer by mananth(16946)

percent percentage ---------------- quantity
Salt solution 10 ---------------- x liters
Salt solution 40 ---------------- 20 - x liters
Mixture 25.00% ---------------- 20

10 x + 40 ( 20 - x ) = 25 * 20

10x+800- 40x=500
10x-40x= 500-800
-30x=-300
/-30
x=10 liters 10.00% Salt solution
10 liters 40.00% Salt solution